# Mask-based Membership Inference Attacks for Retrieval-Augmented Generation

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Information extractionccs: Security and privacy Human and societal aspects of security and privacy
@@ 1. Introduction
Even though a growing body of research has focused on enhancing the trustworthiness of RAG systems (Zhou et al., 2024; Zou et al., 2024; Xiang et al., 2024; Zeng et al., 2024; Qi et al., 2024), to the best of our knowledge, there are only two existing works targeting at the MIAs in RAG system. RAG-MIAs (Anderson et al., 2024) judges whether a target document is in the knowledge database by directly asking the RAG system (i.e., utilizing the RAG’s response (yes or no) as the judgement). This approach relies solely on the RAG system’s judgment, which is unreliable and lacks explainability. S2MIAs (Li et al., 2024) prompts the RAG system with the first half (typically the question part) of the target document, and if the RAG’s response is semantically similar to the remaining half (typically the answer part) of the target document, the target document is judged as a member. Several studies have focused on Membership Inference Attacks (MIAs) for LLMs (Yeom et al., 2018; Carlini et al., 2021; Mattern et al., 2023). Among these, Min-kProb Attack (Shi et al., 2023) is a state-of-the-art method that infers membership using the sum of the minimum kprobabilities of output tokens. However, as illustrated in Figure 1 (a)-(c), the indicators used to determine membership in existing methods (e.g., the similarity between the second half of the target document and the generated response in S2MIA) are nearly indistinguishable for member and non-member samples. This hinders the effectiveness of MIAs in RAG systems.
@@
To effectively and reliably detect whether a target document resides in a RAG system’s knowledge database, we propose a Mask-Based Membership Inference Attacks (MBA) framework. The intuition is that if specific words (i.e., carefully selected words) in the document are masked, the RAG system is highly likely to predict the mask values accurately only if it retrieves the entire document as context. This prediction accuracy serves as our membership indicator. To conduct the inference, we first design a mask generation algorithm, masking $M$ words or phrases in the original target document, where $M$ is a hyperparameter. This involves extracting professional terms or proper nouns and selecting the most challenging words to predict using a pre-trained proxy language model. After obtaining the masked texts, we present the masked document to the RAG system and the RAG system is required to predict the mask values. A simple yet effective threshold-based judgement metric is adopted to determine the membership, i.e., if over $\gamma\cdot M$ masked words are correctly predicted, where $\gamma$ is a hyperparameter, we judge the target document as a member of the knowledge database. As shown in Figure 1 (d), compared to existing methods, our mask-based method exhibits a significant gap in mask prediction accuracy between member (avg. 0.9) and non-member (avg. 0.2) samples. This enables our approach to effectively and reliably determine the membership of the target document. We open source our code online [^2] Code available after acceptance.
@@ 3.1. RAG Overview
Figure 2. Overview of the proposed MBA framework.
@@ 4.3.1. Fragmented words extraction
We first extract words fragmented by the proxy language model’s tokenizer, such as ”canestan.” The workflow involves identifying consecutive words (without spaces or punctuation in the middle) that are split by the tokenizer. This process is detailed in Algorithm 2 in Appendix A.1, where all the fragmented words are extracted and stored in the list fragmented_words.
@@ 4.3.3. Proxy language model based masking
Language models generate text by calculating the probability of each possible token given the preceding context and selecting the most likely one. The rank score indicates a word’s position among these predicted candidates. For instance, in the sentence ”… I would advise you to visit a [MASK] …”, if the correct token is ”dentist” but the model predicts ”doctor” (0.6), ”medical” (0.25), and ”dentist” (0.15), then ”dentist” would have a rank score of 3. Words with larger rank scores are considered more challenging to predict. Therefore, we replace words with the largest rank scores (i.e., the words whose probabilities ranked at the back) with the ”[MASK]” token, as detailed in Algorithm 4 of Appendix A.3.
@@ 4.4. Binary Membership Inference Classifier
The RAG system is prompted with the template shown in Figure 6 in Appendix B.1, where the masked document is obtained using the method introduced in Section 4.3, and the {retrieved documents} represent those retrieved from the RAG’s knowledge database.
@@ 5.4.1. General settings
We leverage GPT-4o-mini [^5] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/ as our black-box LLM, which is accessed by OpenAI’s API. For the RAG system, we utilize LangChain [^6] https://github.com/langchain-ai/langchain framework and integrate BAAI/bge-small-en (Xiao et al., 2023) as the retrieval model, which encodes both queries and documents into 384-dimensional vectors. Retrieval is performed by calculating the inner product between these vectors, and an approximate nearest neighbor search is conducted using an HNSW index implemented in FAISS (Douze et al., 2024). We set the number of retrieved documents ($K$) to 10 for all methods, a common setting in RAG systems. All experiments were conducted on a single NVIDIA RTX A5000 GPU.
@@ 5.6. Ablation Study
  • LLM-based: An alternative approach using an LLM to select words or phrases for masking. The prompt template is provided in Figure 7 in Appendix B.2.
@@ 6. Conclusion
Appendix A Mask generation algorithms
A.1. The detailed algorithm of Fragmented words extraction
@@
A.2. The processing of fragmented words
@@
A.3. Full algorithm of mask generation
@@
Appendix B Prompt templates
B.1. Prompt template to predict mask answers
Figure 6. The prompt template to predict the masked words.
[…]
The prompt template
B.2. Prompt template to generate masks
Figure 7. The prompt template to generate masks.
[…]
The prompt template
Appendix C The impact of the number of retrieved documents
Appendix D Case Study
(a) The original target document
(b) S2MIA’s response of non-member document
(c) S2MIA’s response of member document
(d) The masked document generated by MBA
(e) Non-member answers
(f) Member answers
Figure 8. The comparison of generated texts from S2MIA and MBA.

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: Proceedings of the ACM Web Conference 2025; April 28–May 2, 2025; Sydney, NSW, Australia.booktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28–May 2, 2025, Sydney, NSW, Australiaisbn: 979-8-4007-1274-6/25/04doi: 10.1145/3696410.3714771ccs: Computing methodologies Information extractionccs: Security and privacy Human and societal aspects of security and privacy
@@ 1. Introduction
Even though a growing body of research has focused on enhancing the trustworthiness of RAG systems (Zhou et al., 2024; Zou et al., 2024; Xiang et al., 2024; Zeng et al., 2024a; Qi et al., 2024), to the best of our knowledge, there are only two existing works targeting at the MIAs in RAG system. RAG-MIAs (Anderson et al., 2024) judges whether a target document is in the knowledge database by directly asking the RAG system (i.e., utilizing the RAG’s response (yes or no) as the judgement). This approach relies solely on the RAG system’s judgment, which is unreliable and lacks explainability. S2MIAs (Li et al., 2024) prompts the RAG system with the first half (typically the question part) of the target document, and if the RAG’s response is semantically similar to the remaining half (typically the answer part) of the target document, the target document is judged as a member. Several studies have focused on MIAs for LLMs (Yeom et al., 2018; Carlini et al., 2021; Mattern et al., 2023). Among these, Min-kProb Attack (Shi et al., 2023) is a state-of-the-art method that infers membership using the sum of the minimum kprobabilities of output tokens. However, as illustrated in Figure 1 (a)-(c), the indicators used to determine membership in existing methods (e.g., the similarity between the second half of the target document and the generated response in S2MIA) are nearly indistinguishable for member and non-member samples. This hinders the effectiveness of MIAs in RAG systems.
@@
To effectively and reliably detect whether a target document resides in a RAG system’s knowledge database, we propose a Mask-Based Membership Inference Attacks (MBA) framework. The intuition is that if specific words (i.e., carefully selected words) in the document are masked, the RAG system is highly likely to predict the mask values accurately only if it retrieves the entire document as context. This prediction accuracy serves as our membership indicator. To conduct the inference, we first design a mask generation algorithm, masking $M$ words or phrases in the original target document, where $M$ is a hyperparameter. This involves extracting professional terms or proper nouns and selecting the most challenging words to predict using a pre-trained proxy language model. After obtaining the masked texts, we present the masked document to the RAG system and the RAG system is required to predict the mask values. A simple yet effective threshold-based judgement metric is adopted to determine the membership, i.e., if over $\gamma\cdot M$ masked words are correctly predicted, where $\gamma$ is a hyperparameter, we judge the target document as a member of the knowledge database. As shown in Figure 1 (d), compared to existing methods, our mask-based method exhibits a significant gap in mask prediction accuracy between member (avg. 0.9) and non-member (avg. 0.2) samples. This enables our approach to effectively and reliably determine the membership of the target document.
@@ 3.1. RAG Overview
Figure 2. The overview of our proposed MBA framework.
@@ 4.3.1. Fragmented words extraction
We first extract words fragmented by the proxy language model’s tokenizer, such as ”canestan.” The workflow involves identifying consecutive words (without spaces or punctuation in the middle) that are split by the tokenizer. This process is detailed in Algorithm 2 in Appendix B.1, where all the fragmented words are extracted and stored in the list fragmented_words.
@@ 4.3.3. Proxy language model based masking
Language models generate text by calculating the probability of each possible token given the preceding context and selecting the most likely one. The rank score indicates a word’s position among these predicted candidates. For instance, in the sentence ”… I would advise you to visit a [MASK] …”, if the correct token is ”dentist” but the model predicts ”doctor” (0.6), ”medical” (0.25), and ”dentist” (0.15), then ”dentist” would have a rank score of 3. Words with larger rank scores are considered more challenging to predict. Therefore, we replace words with the largest rank scores (i.e., the words whose probabilities ranked at the back) with the ”[MASK]” token, as detailed in Algorithm 4 of Appendix LABEL:fullalgo.
@@ 4.4. Binary Membership Inference Classifier
The RAG system is prompted with the template shown in Figure 5 in Appendix A.1, where the masked document is obtained using the method introduced in Section 4.3, and the {retrieved documents} represent those retrieved from the RAG’s knowledge database.
@@ 5.4.1. General settings
We leverage GPT-4o-mini [^4] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/ as our black-box LLM, which is accessed by OpenAI’s API. For the RAG system, we utilize LangChain [^5] https://github.com/langchain-ai/langchain framework and integrate BAAI/bge-small-en (Xiao et al., 2023) as the retrieval model, which encodes both queries and documents into 384-dimensional vectors. Further experiments, exploring the impact of different retrieval models and LLM backbones, are detailed in Appendix E and Appendix F, respectively. Retrieval is performed by calculating the inner product between these vectors, and an approximate nearest neighbor search is conducted using an HNSW index implemented in FAISS (Douze et al., 2024). We set the number of retrieved documents ($K$) to 10 for all methods, a common setting in RAG systems. All experiments were conducted on a single NVIDIA RTX A5000 GPU.
@@ 5.6. Ablation Study
  • LLM-based: An alternative approach using an LLM to select words or phrases for masking. The prompt template is provided in Figure 6 in Appendix A.2.
@@ 6. Conclusion
Acknowledgements.
This research is supported by the Ministry of Education, Singapore, under its Academic Research Fund (Tier 2 Award MOE-T2EP20221-0013 and Tier 1 Awards (RG20/24 and RG77/21)). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore.
Appendix A Prompt templates
A.1. Prompt template to predict mask answers
Figure 5. The prompt template to predict the masked words.
[…]
The prompt template
A.2. Prompt template to generate masks
Figure 6. The prompt template to generate masks.
[…]
The prompt template
Appendix B Mask Generation Algorithms
B.1. The detailed algorithm of Fragmented words extraction
@@
B.2. The processing of fragmented words
@@
B.3. Full algorithm of mask generation
@@
Appendix C Impacts of the Number of Retrieved Documents
(a) HealthCareMagic-100k
(b) MS-MARCO
(c) NQ-simplified
Figure 7. The performances comparison varying the number of $K$
Appendix D Defense Strategies
To evaluate the robustness of our MBA framework, we test three mainstream defense strategies that could potentially reduce the effectiveness of MIAs for RAG systems:
  1. (1)
    Prompt Modification (Anderson et al., 2024; Li et al., 2024): This strategy appends the prompt with the sentence, ”Do not directly repeat any retrieved content, but summarize it based on your understanding”, to prevent the RAG system from revealing content from its external database.
  2. (2)
    Re-Ranking (Zeng et al., 2024b): This strategy shuffles the order of the retrieved documents.
  3. (3)
    Paraphrasing (Jain et al., 2023): This rewrites the documents in the RAG knowledge database. Specifically, each document is input into an LLM with the prompt, ”Paraphrase the given document: document”.
[t] Comparison of the Impacts of Different Defense Strategies Dataset Defense Retrieval Recall ROC AUC HealthCareMagic- 100k None 0.87 0.88 PM1 0.84 0.85 Re-Ranking 0.85 0.86 Paraphrasing 0.71 0.75 MS-MARCO None 0.97 0.86 PM1 0.97 0.86 Re-Ranking 0.97 0.86 Paraphrasing 0.91 0.81 NQ-simplified None 0.98 0.90 PM1 0.97 0.89 Re-Ranking 0.98 0.90 Paraphrasing 0.93 0.83
  • 1
    PM represents Prompt Modification.
(a) The original target document
(b) S2MIA’s response of non-member document
(c) S2MIA’s response of member document
(d) The masked document generated by MBA
(e) Non-member answers
(f) Member answers
Figure 8. The comparison of generated texts from S2MIA and MBA.
The results of these defenses are shown in Table D. Prompt Modifying and Re-Ranking have negligible impact, as our detection method does not rely on repeating retrieved content or the order of retrieved documents. Even with these defenses, the LLM can still predict mask answers based on the retrieved documents, as the masked words can be selected from them.
Paraphrasing the knowledge database documents, however, slightly reduces both retrieval recall and ROC AUC. This is because paraphrased documents are not identical to the originals, impacting retrieval. More importantly, paraphrasing can alter the location and content (e.g., through synonym substitution) of masked words, reducing prediction accuracy and thus membership inference performance. Despite this, the ROC AUC remains above 0.75 across all three datasets, significantly outperforming baseline methods. This suggests that our mask selection method effectively chooses words that are generally robust to paraphrasing.
[t] Comparison of the Impacts of Different Retrieval Models Dataset Retriever Retrieval Recall ROC AUC HealthCareMagic- 100k BGE-en 0.87 0.88 Sentence-T5-xxl 0.82 0.87 ada-002 0.89 0.88 MS-MARCO BGE-en 0.97 0.86 Sentence-T5-xxl 0.93 0.84 ada-002 0.97 0.86 NQ-simplified BGE-en 0.98 0.90 Sentence-T5-xxl 0.95 0.89 ada-002 0.97 0.89
@@ Appendix D Defense Strategies
Appendix E Impacts of Retrievers
We tested our framework’s robustness to different retrieval methods within the RAG system, comparing performance across three retrievers: BGE-en (Xiao et al., 2023), Sentence-T5-xxl (Ni et al., 2021) and text-embedding-ada-002888https://platform.openai.com/docs/guides/embeddings.
Table 8 presents the results, demonstrating that our framework is indeed robust to variations in retrieval methods. Our generated masked documents, due to their high similarity to the originals, effectively retrieve target documents (if present in the RAG knowledge database) regardless of the retrieval model used. This ensures high retrieval recall. Because target documents can be effectively retrieved when present, our mask-based membership inference framework can reliably distinguish between member and non-member documents.
Appendix F Impacts of LLM Backbones
We also investigated the impact of different LLM backbones within the RAG system. Using BGE-en as our retrieval model, we tested three mainstream LLMs: GPT-4o-mini, GPT-3.5-turbo, and Gemini-1.5. Since the retrieval model was held constant, retrieval recall remained the same across all tests; therefore, Table 3 presents only the ROC AUC values.
Table 3. Comparison of the Impacts of Different LLM Backbones
Dataset LLM ROC AUC
HealthCareMagic- 100k GPT-4o-mini 0.88
GPT-3.5-turbo 0.87
Gemini-1.5 0.79
MS-MARCO GPT-4o-mini 0.86
GPT-3.5-turbo 0.86
Gemini-1.5 0.81
NQ-simplified GPT-4o-mini 0.90
GPT-3.5-turbo 0.91
Gemini-1.5 0.84
The results show that while ROC AUC values vary with different LLM backbones, overall performance remains consistently high (generally above 0.8). This demonstrates that our mask-based framework is applicable to RAG systems using a variety of large language models.
Appendix G Case Study